# Predicting the Performance of Black-box LLMs through Self-Queries

### v2

@@ Abstract
As large language models (LLMs) are increasingly relied on in AI systems, predicting when they make mistakes is crucial. While a great deal of work in the field uses internal representations to interpret model behavior, these representations are inaccessible when given solely black-box access through an API. In this paper, we extract features of LLMs in a black-box manner by using follow-up prompts and taking the probabilities of different responses as representations to train reliable predictors of model behavior. We demonstrate that training a linear model on these low-dimensional representations produces reliable and generalizable predictors of model performance at the instance level (e.g., if a particular generation correctly answers a question). Remarkably, these can often outperform white-box linear predictors that operate over a model’s hidden state or the full distribution over its vocabulary. In addition, we demonstrate that these extracted features can be used to evaluate more nuanced aspects of a language model’s state. For instance, they can be used to distinguish between a clean version of GPT-4o-mini and a version that has been influenced via an adversarial system prompt that answers question-answering tasks incorrectly or introduces bugs into generated code. Furthermore, they can reliably distinguish between different model architectures and sizes, enabling the detection of misrepresented models provided through an API (e.g., identifying if GPT-3.5 is supplied instead of GPT-4o-mini).
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong performance on a wide variety of tasks (Radford et al., 2019), leading to their increased involvement in larger systems. For instance, they are often used to provide supervision (Bai et al., 2022; Sam et al., 2024), as tools in decision-making (Benary et al., 2023; Sha et al., 2023), or as controllers on agentic frameworks (Xi et al., 2023; Robey et al., 2024). Thus, it is crucial to understand and predict their behaviors, especially in high-stakes settings. However, as with any deep network, it is difficult to understand the behavior of such large models (Zhang et al., 2021). For instance, prior work has studied input gradients or saliency maps (Simonyan et al., 2013; Zeiler and Fergus, 2014; Pukdee et al., 2024)) to attempt to understand neural network behavior, but this can fail to reliably describe model behavior (Adebayo et al., 2018; Kindermans et al., 2019; Srinivas and Fleuret, 2020). Other work has studied the ability of transformers to represent certain algorithms (Nanda et al., 2022; Zhong et al., 2024) that may be involved in their predictions.
@@
One promising direction in understanding LLMs (or any other multimodal model that understands natural language) is to leverage their ability to interact with human queries. Recent work has demonstrated that a LLM’s hidden state contains low-dimensional features of model truthfulness or harmfulness (Zou et al., 2023a). Other work studies learning sparse dictionaries and analyzing how these networks activate on certain, related input tokens (Bricken et al., 2023). While significant progress has been made on these fronts, these approaches all require white-box access to these models (i.e., access to the model’s activations or hidden states). However, many of the best-performing LLMs (Achiam et al., 2023; Team et al., 2023) lie beyond closed-source APIs, so these prior attempts to understand model behavior cannot be applied. This raises the question, “How well can we predict the LLM’s behavior with only black-box access?
Figure 1: Our approach to extract black-box representations from LLMs which can be used for various applications, including predicting performance, determining if correct models are given through an API, and detecting models that have been influenced by adversarial system prompts.
In this paper, we propose to extract useful representations for predicting the performance of black-box LLMs by querying them about their outputs. In essence, after receiving a generation or prediction from a LLM, we leverage the LLM’s ability to reason about its own generated answer and meaningfully respond to follow-up questions, such as, “Are you able to explain your answer?” Our hypothesis is that the probability distribution over answers to these questions significantly varies between whether the model’s original answer is correct, as well as for different model classes and sizes. As we only look at the outputs of these LLMs (i.e., top-k token probabilities that are accessible through many APIs), we remark that this approach is both model-agnostic and works for closed-source models. When top-k probabilities are not provided, we can approximate this by sampling from the LLM, and we provide a result on how quickly this approximation converges to the approach with the true underlying probabilities under the LLM.
In our experiments, we demonstrate that our proposed approach of querying a model with elicitation questions produces features useful in various applications. We first demonstrate that they can be used to train accurate predictors of model performance (e.g., predicting whether a particular class prediction or text generation is correct). Our approach of querying a model with elicitation questions often matches or outperforms linear predictors that operate over the LLM’s hidden state (i.e., requiring white-box access), over a wide variety of LLMs applied to question-answering (QA) tasks. As our extracted features are low-dimensional, we also observe that predictors trained on them have stronger generalization guarantees. We also observe that sampling-based approximations closely match the performance of a model that uses the true probabilities, so our method performs well even without access to top-k probabilities. Finally, we study the role of diversity in these questions, with the interesting finding that even diverse unrelated sequences of natural language (i.e, not in the form of questions) can sometimes outperform using specific elicitation questions.
In addition to predicting LLM performance at the example level, these extracted representations are also useful for a variety of other applications in assessing the state of a LLM. For instance, recent work demonstrated that model internals can be used to assess when an LLM has been adversarially influenced by a prompt (MacDiarmid et al., 2024) to exhibit harmful behavior. Our work extends this setting by demonstrating that our extracted representations can be used to almost perfectly detect when a LLM (e.g., GPT-3.5) has been adversarially influenced by a system prompt in a completely black-box fashion. We also provide evidence that our approach is robust to variations in the system prompt through various ablation studies. Finally, we also demonstrate that our approach can be used to reliably distinguish between different model architectures and model sizes; this can be useful in evaluating if cheaper or smaller models are falsely being provided through these closed-source APIs. Overall, our work demonstrates that follow-up queries and interactions with a black-box LLM API can produce useful features in predicting its behavior — providing promising results that can sometimes even match the performance of looking at activations.
@@ 2 Related Work
Predicting the behavior of deep neural networks is an important problem in the field, due to the difficult-to-interpret nature of these models. Existing work looks to assess the performance of models by directly operating over the weight space (Unterthiner et al., 2020) or ensembles of multiple trained models (Jiang et al., 2021). Specifically for language models, prior work has primarily focused on predicting task-level performance on new tasks; for instance, developing predictors of task-level performance that use the performance on similar or related tasks (Xia et al., 2020; Ye et al., 2023). Other work attempts to predict the performance of models as they scale up computation (often in terms of data and model size) (Kaplan et al., 2020; Muennighoff et al., 2024). Our work is different as we predict instance-level performance (i.e., correctness on a certain input), and we leverage a small amount of labeled data from the downstream task.
@@
Many other works have explored approaches to extract representations from neural networks (NNs). A related line of work looks to train NNs (specifically image classifiers) to extract a small set of discrete, interpretable concepts, which can be passed through a linear probe to recover a classifier (Koh et al., 2020). In our case, we leverage the ability of the LLM to understand language and can circumvent this need for training, extracting features in a task-agnostic manner. Prior work has studied how to extract useful representations for downstream tasks (Wang et al., 2023; Zou et al., 2023a). Our approach significantly differs in nature from these approaches, as we are looking to extract more compressed, low-dimensional features that reveal information about black-box model behavior. Perhaps the most related work employs a similar strategy of asking questions, specifically to detect instances where a model is untruthful (Pacchiardi et al., 2024). Our work encompasses the broader task of predicting model behavior and performance.
Uncertainty Quantification in LLMs
Finally, a related line of work is assessing the calibration or ability of a language model to represent its own uncertainty (Xiong et al., 2023). Many of the follow-up, elicitation questions that we ask prompt the model to look at its answer and answer “Yes” or “No”; this is related to the notion of a model’s ability to understand what it knows (Kadavath et al., 2022; Kapoor et al., 2024) or reflect uncertainty in its own decisions. While we primarily access top-$k$ probabilities in its responses to construct our representations, existing work using verbalized confidence scores One related work looks at iterative prompts as a better way to extract different notions of uncertainty (epistemic vs. aleatoric) (Yadkori et al., 2024). Our work is different from this line of work, however, as we elicit these probabilities as a representation from such a model to train a simple, calibrated linear classifier, for predicting model performance and for auditing adversarially influenced LLMs or different model architectures. We provide a comparison to a variety of uncertainty quantification methods, empirically showing many benefits of extracting additional information with our approach of using follow-up queries.
3 Eliciting Black-box Representations from Language Models
@@
As we do not assume access to the internals of a LLM, we propose to extract useful features in predicting its behavior by asking follow-up questions about its generations. This is completely black-box as we only look at the model’s outputs, or more specifically, its top-$k$ probabilities over the next token. We feed these as features into simple linear classifiers for some downstream task (e.g., predicting performance). For some APIs, we do not have access to the LLM’s top-$k$ probabilities, so we theoretically analyze predictors trained on sampled approximations of these probabilities.
3.1 Extracting Features by Follow-up Questions
@@
To extract our black-box representations, we prompt the model with a large number of elicitation questions. We consider a set of questions $Q=\{q_{1},...,q_{d}\}$ and some autoregressive language model, which models some distribution $P$ over sequences of text. We also consider a dataset $D=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$, where $x_{i}$ is a sequence of tokens and $y_{i}$ corresponds to a binary label, for example, if the LLM has correctly answered the question $x_{i}$. We define $a_{i}$ as the greedy response from the LLM, or that $a_{i}=\arg\max_{c}P(c|x_{i})$. Then, we construct our black-box representation as some vector $z=(z_{1},...,z_{d})$, where each $z_{j}=P(\texttt{yes}|x\oplus a\oplus q_{j})$, where $\oplus$ denotes concatenation. In other words, dimensions of our representation correspond to the probability of the yes token under the LLM (where the distribution is specified over the yes and no tokens), in response to the question $x$, the greedy sampled answer $a$, and the elicitation question $q_{j}$. The elicitation questions are detailed in Appendix D.4, but generally consist of simple self-inquiry questions such as “Do you think your answer is correct?” or “Are your responses free from bias?” This simple approach allows us to add more information to representations by continuing to generate new follow-up questions. In our paper, we find that working with a set of roughly 50 questions seems to be sufficient for strong performance (see ablations in Section 4.5).
@@
In addition to these probabilities of responses to questions, we also append: (1) pre- and post-confidence scores of the LLM, which are responses to asking the question before and after generating a greedy sample from the model, and (2) the distribution over possible answers for the task, (for open-ended QA tasks, we simply use the log probability of the greedy output). In our experiments with GPT models, we append the sorted top-5 probabilities returned by the API. We train a linear predictor $\beta$ to predict the label $y$ (e.g., whether the model is correct or not) given our feature vector $z$.
3.2 Generating Follow-up Prompts
To construct this set of eliciting questions $Q$, we specify a small number of questions that relate to the model’s confidence or belief in its answer. We also use GPT4 to generate a larger number (40) of questions. The questions and prompts used to generate the GPT4-generated questions are given in Appendix D.4. As noted in prior work that uses similar questions for lie detection (Pacchiardi et al., 2024), a wide variety of questions seems to lead to more useful representations, capturing more information from the LLM than say directly querying it about its answer (e.g., standard approaches in uncertainty quantification).
@@ 3.2 Generating Follow-up Prompts
We note that based on the specific nature of the question, the response (e.g., the probability of responding yes) could define a weak predictor of whether the model is correct or not. This is reminiscent of the design of weak learners in boosting (Freund and Schapire, 1996) or weak labelers in programmatic weak supervision (Ratner et al., 2017; Sam and Kolter, 2023; Smith et al., 2024). However, to maintain our approach’s generality and to not restrict our approach to only a certain type of elicitation questions, we treat these as abstract features for a linear predictor. We also note that further work could perform discrete optimization over prompts to further improve the extracted representation’s usability, through methods described in (Wen et al., 2024; Zou et al., 2023b; Chao et al., 2023). However, one key appeal of the current approach is that it defines an extremely simple classifier in a task-agnostic fashion. Performing optimization over these questions might lead to overfitting, and the resulting predictors on the outputs of these prompts require more complex analysis in deriving valid generalization bounds.
3.3 Analysis on Finite Samples from Black-box LLMs
@@
While our approach described above assumes access to the top-$k$ probabilities, some LLMs are only accessible through APIs that do not provide this information (Team et al., 2023). In this setting, we can approximately compute these probabilities via high-temperature sampling from the LLM. Here, we provide a theoretical analysis of how this approximation impacts the performance of our method.
@@
Recall that we have our representation $z=(z_{1},...,z_{d})$, which corresponds to the actual probability of the yes token under the LLM. Without access to these true probabilities through an API, we instead have some approximation $\hat{z}=(\hat{z}_{1},...,\hat{z}_{d})$, where each $\hat{z}_{j}$ is an average of $k$ samples from $\text{Ber}(z_{j})$. From prior work in logistic regression under settings of covariate measurement error (Stefanski and Carroll, 1985), when we have that $k$ grows with $n$, we observe that the naive MLE (maximum likelihood estimator) on the observed approximation results in a consistent, albeit biased, estimator. We present an analysis of our setting, showing a result on the convergence rate of the MLE for $\beta$.
@@
Let $\hat{\beta}$ be the MLE for the logistic regression on the dataset $\{(x_{i}^{j},y_{i})|i=1,...,n,j=1,...,k\}$, where $x_{i}^{j}$ are independent samples from Ber$(p_{i})$. We assume there exists some unique optimal set of weights $\beta_{0}$ over inputs $p=(p_{1},...,p_{d})$, and we let $n,k>>d$. Then, we have that $\hat{\beta}\to\beta_{0}$ as $n\to\infty$ and $k\to\infty$. Furthermore, $\hat{\beta}$ converges at a rate $O\left(\frac{1}{\sqrt{n}}+\frac{\sqrt{n}}{k}\right)$.
We provide the proof of this statement in Appendix B. At a high level, this follows from relatively standard results; $\hat{\beta}$ converges to the optimal predictor on the sampled dataset (which we call $\beta^{*}$) via asymptotic results for the MLE. Then, we derive that $\beta^{*}$ converges to $\beta_{0}$ at a rate of $O\left(\sqrt{n}/k\right)$.
This result demonstrates that, under the setting where we do not have access to the LLM’s actual probabilities, we can closely approximate this with sampling, as long as we approximate it with a sample of size $k$ that grows (at a slower rate) with $n$ to get a consistent estimator. Later in Section 4.5, we empirically demonstrate that a naive logistic regression model with an approximation over a finite $k$ samples performs comparably to using the actual LLM probabilities.
Figure 2: AUROC in predicting model performance on the open-ended QA benchmarks of Natural Questions (Top) and SQuAD (Bottom). Dashed bars represent white-box methods, which assume more access than QueRE. RepE cannot be applied to black-box models (e.g., GPT models). Full Logits for the GPT models is an approximation of a sparse vector with nonzero values for the top-5 logits from the API.
Figure 3: AUROC in predicting model performance on closed-ended QA benchmarks of HaluEval, BoolQ, and DHate. Dashed bars represent white-box methods. RepE cannot be applied to black-box models (e.g., GPT models). Full Logits for GPT models is a sparse vector with nonzero values for the top-5 logits from the API.

### v4

# Predicting the Performance of Black-box Language Models with Follow-up Queries

@@ Abstract
Reliably predicting the behavior of language models—such as whether their outputs are correct or have been adversarially manipulated—is a fundamentally challenging task. This is often made even more difficult as frontier language models are offered only through closed-source APIs, providing only black-box access. In this paper, we predict the behavior of black-box language models by asking follow-up questions and taking the probabilities of responses as representations to train reliable predictors. We first demonstrate that training a linear model on these responses reliably and accurately predicts model correctness on question-answering and reasoning benchmarks. Surprisingly, this can even outperform white-box linear predictors that operate over model internals or activations. Furthermore, we demonstrate that these follow-up question responses can reliably distinguish between a clean version of an LLM and one that has been adversarially influenced via a system prompt to answer questions incorrectly or to introduce bugs into generated code. Finally, we show that they can also be used to differentiate between black-box LLMs, enabling the detection of misrepresented models provided through an API. Overall, our work shows promise in monitoring black-box language model behavior, supporting their deployment in larger, autonomous systems.
@@ 1 Introduction
Reliably predicting the behavior of a language model (e.g., whether its outputs are correct, or whether it has been adversarially manipulated) is a fundamentally challenging task. This is made even more challenging as many of the most capable large language models (LLMs) lie beyond closed-source APIs (achiam2023gpt; team2023gemini), providing only black-box access through inputs and outputs. As a result, recent advances in understanding these models through model internals or from mechanistic viewpoints (olsson2022context; nanda2022progress) are no longer applicable. The inability to rely on LLMs remains a roadblock for their widespread deployment in high-stakes settings or in agentic and autonomous frameworks (xi2023rise; robey2024jailbreaking).
@@
In spite of only having black-box access, a promising direction in understanding LLMs is to leverage their ability to interact with human queries and provide useful responses. Recent work in the white-box setting (i.e., having access to model internals) has demonstrated that a language model’s hidden state contains low-dimensional features of truthfulness or harmfulness (zou2023representation), and has analyzed learning sparse dictionaries and activations on certain input tokens (bricken2023monosemanticity). While significant progress has been made on these fronts, these approaches all require white-box access to these models. This raises the question, “How well can we predict a language model’s behavior with only black-box access?"
Figure 1: Our approach predicts LLM behavior using linear predictors trained on features derived from follow-up questions posed to the LLM. We show that responses to follow-up questions are highly predictive of correctness on downstream benchmarks, and are useful in distinguishing between black-box models and for detecting if models have been influenced by an adversary.
In this paper, we propose to predict model behavior by looking at their responses to follow-up questions. After receiving an initial generation or answer from an LLM, we ask a set of follow-up questions, such as, “Are you able to explain your answer?” We then take the probability of the ‘‘Yes’’ token of its response as our features for predicting model behavior. Our hypothesis is that the distributions over answers to these questions meaningfully vary with correctness, model families, and model scale. A key advantage of our approach is that, because it only relies on model outputs, it is also model-agnostic and broadly applicable. In cases where top-k probabilities are not available, we can approximate them via sampling. We provide a theoretical result on how quickly using this approximation converges to the approach that has the true underlying probabilities from the LLM.
Our experiments demonstrate that querying a model with follow-up questions yields features that are highly predictive of performance on LLM benchmarks. We show that simple linear models trained on these features accurately predict instance-level correctness on question-answering and reasoning tasks. Surprisingly, our black-box approach often matches—or even outperforms—white-box methods that operate over the language model’s hidden state, across a range of different language models and benchmarks. Furthermore, we demonstrate that our predictors admit nice generalization properties due to their low-dimensional nature and perform well on out-of-distribution data (e.g., transferred to new model scales or new datasets) due to our approach’s generality.
Beyond predicting performance on benchmarks, our approach provides insights into other model behaviors. For instance, these follow-up questions can be used to reliably detect when an LLM (e.g., GPT-4o-mini) has been adversarially influenced via a system prompt to generate incorrect answers or introduce hidden bugs into code. We also demonstrate that these follow-up question responses can be used to accurately distinguish between different black-box LLMs; this is useful in auditing if cheaper or smaller models are falsely being provided through closed-source APIs. Together, these results highlight the promise of our approach in predicting and monitoring the behaviors of black-box language models, supporting their future use in large systems.
@@ 2 Related Work
As previously mentioned, predicting the performance deep learning models is challenging due to their difficult-to-interpret nature. Existing work looks to assess the performance of models by directly operating over the weight space (unterthiner2020predicting) or ensembles of multiple trained models (jiang2021assessing). Specifically for language models, prior work has primarily focused on predicting task-level performance on new tasks; for instance, developing predictors of task-level performance that use the performance on similar or related tasks (xia2020predicting; ye2023predictable). Other work attempts to predict the performance of models as they scale up computation (often in terms of data and model size) (kaplan2020scaling; muennighoff2024scaling). In contrast, our work focuses on instance-level prediction—i.e., determining whether a model’s response to a specific input is likely to be correct. Furthermore, we operate in a black-box setting, using only input-output behavior, rather than internal model parameters or activations.
Uncertainty Quantification in LLMs
A related line of work is assessing the calibration or ability of a language model to represent its own uncertainty (xiong2023can). Some work investigates LLMs’ ability to verbalize confidence or self-assess the quality of their outputs (kadavath2022language; kapoor2024large), and others explore prompting techniques to elicit richer uncertainty estimates—e.g., distinguishing between epistemic and aleatoric uncertainty via iterative queries (yadkori2024believe). Our approach is related in that we ask follow-up questions (e.g., “Are you confident in your answer?”) to elicit indicators of model uncertainty. However, we differ in our use of these responses: rather than relying on a model’s verbalized confidence alone, we extract token-level probabilities as features and train simple linear classifiers to predict correctness. We further show that these features generalize across models and settings, and are useful for a large set of tasks that go beyond the set of calibration metrics focused on in the uncertainty quantification literature. In fact, we provide a comparison with a variety of uncertainty quantification methods, empirically showing many benefits of extracting additional information with multiple follow-up queries.
@@
Many other works have explored approaches to extract representations from neural networks. A related line of work looks to train neural networks (specifically image classifiers) to extract a small set of discrete, interpretable concepts, which can be passed through a linear probe to recover a classifier (koh2020concept). In our case, we leverage the ability of the LLM to understand language and can circumvent this need for training, extracting features in a task-agnostic manner. Prior work has studied how to extract useful representations for downstream tasks (wang2023improving; zou2023representation), although they operate in the fundamentally different white-box setting where you can access model internals. Perhaps the most related work employs a similar strategy of asking questions, specifically to detect instances where a model is untruthful (pacchiardi2024how). Our work encompasses the much broader task of predicting model behavior and performance.
3 Predicting Performance with Follow-up Queries
@@
Without any access to language model internals, we propose to elicit useful features about its behavior by asking follow-up questions about its generations. This is completely black-box as we only look at the model’s outputs, or more specifically, its top-$k$ probabilities over the next token. We feed these as features into simple linear classifiers for some downstream task (e.g., predicting performance). For some APIs, we do not have access to the LLM’s top-$k$ probabilities, so we theoretically analyze predictors trained on sampled approximations of these probabilities.
3.1 Predictive Features through Follow-up Responses
@@
We consider a set of follow-up queries $Q=\{q_{1},...,q_{d}\}$ and some autoregressive language model, which models some distribution $P$ over sequences of text. We also consider a dataset $D=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$, where $x_{i}$ is a sequence of tokens and $y_{i}$ corresponds to a binary label, for example, if the LLM has correctly answered the question $x_{i}$. We define $a_{i}$ as the greedily sampled response from the LLM, or that $a_{i}=\arg\max_{c}P(c|x_{i})$. Then, we construct our black-box representation as some vector $z=(z_{1},...,z_{d})$, where each $z_{j}=P(\texttt{yes}|x\oplus a\oplus q_{j})$, where $\oplus$ denotes the concatenation of strings (or tokens). Each dimension of our representation corresponds to the probability of the yes token under the LLM (where the distribution is specified over the yes and no tokens), in response to the follow-up question $q_{j}$ about the pair of the original question $x$ and greedily sampled answer $a$. In our paper, we find that working with a set of roughly 50 questions seems to be sufficient for strong performance (see ablations in Section˜4.5). We also analyze different choices of these questions in Appendix A.5. Notably, all features $z$ can be extracted in parallel, so increasing the number of follow-up questions adds minimal computational overhead.
@@
In addition to these features, on closed-ended QA tasks, we append the distribution over possible answers. On both closed-ended and open-ended QA tasks, we append the pre- and post-confidence score, which is the confidence of the language model before and after it sees its own sampled answer. We train a linear predictor $\beta$ to predict the label $y$ (e.g., whether the model is correct or not) given our feature vector $z$.
Generating Follow-up Prompts
To construct this set of eliciting questions $Q$, we specify a small number of questions that relate to the model’s confidence or belief in its answer. We also use GPT4 to generate a larger number (40) of questions. The questions and prompts used to generate the GPT4-generated questions are given in Appendix D.4. The elicitation questions are detailed in Appendix D.4, but generally consist of simple self-inquiry questions such as “Do you think your answer is correct?” or “Are your responses free from bias?” This simple approach allows us to add more information to our extracted representations by continuing to generate new follow-up questions.
@@
We note that, based on the specific nature of the question, the response (e.g., the probability of responding yes) could define a weak predictor of whether the model is correct or not. This is reminiscent of the design of weak learners in boosting (freund1996experiments) or weak labelers in programmatic weak supervision (ratner2017snorkel; sam2023losses; smith2024language). However, to maintain our approach’s generality and to not restrict our approach to only a certain type of elicitation questions, we treat these as abstract features for a linear predictor. We also note that further work could perform discrete optimization over prompts to further improve the extracted representation’s usability, through methods described in (wen2024hard; zou2023universal). However, one key appeal of the current approach is that it defines an extremely simple classifier in a task-agnostic fashion. Performing optimization over these questions might lead to overfitting, and the resulting predictors on the outputs of these prompts require more complex analysis in deriving valid generalization bounds.
3.2 Theoretical Analysis of Sampling-based Approximations
@@
While our approach described above assumes access to the top-$k$ probabilities, some language models are only accessible through APIs that do not provide this information (team2023gemini). In this setting, we can approximate these probabilities via high-temperature sampling from the LLM. Here, we provide a theoretical analysis of how this approximation impacts the performance of our method.
@@
Recall that we have our representation $z=(z_{1},...,z_{d})$, which corresponds to the actual probability of the yes token under the LLM. Without access to these true probabilities through an API, we instead have some approximation $\hat{z}=(\hat{z}_{1},...,\hat{z}_{d})$, where each $\hat{z}_{j}$ is an average of $k$ samples from $\text{Ber}(z_{j})$. From prior work in logistic regression under settings of covariate measurement error (stefanski1985covariate), when we have that $k$ grows with $n$, we observe that the naive MLE (maximum likelihood estimator) on the observed approximation results in a consistent, albeit biased, estimator. We present an analysis of our setting, showing a result on the convergence rate of the MLE for $\beta$.
@@
[…]
We provide the full proof in Appendix B. At a high level, this follows straightforwardly; $\hat{\beta}$ converges to the optimal predictor on the sampled dataset (which we call $\beta^{*}$), via asymptotic results for the MLE. Then, we derive that $\beta^{*}$ converges to $\beta_{0}$ at a rate of $O\left(\sqrt{n}/k\right)$.
This result demonstrates that, under the setting where we do not have access to the LLM’s actual probabilities, we can closely approximate this with sampling, as long as we approximate it with a sample of size $k$ that grows (at a slower rate) with $n$ to get a consistent estimator. Later in Section˜4.5, we empirically demonstrate that a naive logistic regression model with an approximation over a finite $k$ samples performs comparably to using the actual LLM probabilities.
Table 1: AUROC in predicting model performance on the reasoning benchmarks of GSM8k and CodeContests. QueRE performs the best in predicting correctness on reasoning tasks.
Dataset LLM Logits Pre-conf Post-conf Self-Cons. Sem. Entropy QueRE
GSM8K GPT-3.5 0.5636 0.5203 0.4534 0.5227 0.7495 0.7748
GPT-4o-mini 0.5463 0.5539 0.5474 0.5012 0.5546 0.7319
Code Contests GPT-3.5 0.6001 0.4812 0.4244 0.5036 0.5346 0.6800
GPT-4o-mini 0.5274 0.5171 0.5218 0.5000 0.5604 0.7924
Figure 2: AUROC in predicting model performance on the open-ended QA benchmarks of Natural Questions (Top) and SQuAD (Bottom). Dashed bars represent white-box methods, which assume more access than QueRE. QueRE often best predicts model performance on open-ended QA tasks, even when compared to white-box methods.
Figure 3: AUROC in predicting model performance on closed-ended QA benchmarks of HaluEval, BoolQ, and DHate. Dashed bars represent white-box methods.